Given that M is a singular matrix, evaluate x where;



6 7 -1



M = 3 x 5



9 11 x





1
Expert's answer
2022-02-14T18:36:25-0500

"M=\begin{pmatrix}
 6& 7 & -1\\
 3 & x & 5\\
 9& 11 & x
\end{pmatrix}"

det M=0

"\begin{vmatrix}
 6& 7 & -1\\
 3 & x & 5\\
 9& 11 & x
\end{vmatrix}=0\\
6x^2-33+315+9x-330-21x=0\\
6x^2-12x-48=0\\
x^2-2x-8=0\\
x_1=-2, x_2=4"